OpenStudy (anonymous):

A textbook weighs 3 kilograms.How many grams is this? A.300 g B.3,000g C.30,000g D.0.003g @shamil98

OpenStudy (anonymous):

B

OpenStudy (mathstudent55):

There are 1000 g in a kg. 1 kg = 1000 g 2 kg = 2000 g 3 kg = ...

OpenStudy (anonymous):

i thought it was C

OpenStudy (anonymous):

No, it's B.

OpenStudy (doc.brown):

1kg = 1000g\[3kg\times\frac{1000g}{kg}=\]\[3\cancel{kg}\times\frac{1000g}{\cancel{kg}}=\]\[3\times1000g=\]